OpenStudy (anonymous):

Which choices are solutions to the following equation? X^2-3x=-5/4 There are multiple choices I’m pretty sure …

OpenStudy (anonymous):

My choices, feel free to go through with me if you want.

OpenStudy (anonymous):

I think its x=.5 I wouldnt quote me though, because I dont know if I did it correctly

OpenStudy (anonymous):

Okay, thank you for taking the time to reply!

OpenStudy (anonymous):

@DarkFlame52 is correct

OpenStudy (anonymous):

YOU HAVE x^2-3x=-5/4 so crossmultiply 4 on otherside it will be 4x^2-12x+5=0 then factorize it it will be 4x^2-10x-2x+5=0 hence, (2x-1)(2x-5) are factors hence x=2.5which matches the option B of your question...